# Probing lepton flavor violating dark matter scenarios via astrophysical photons and positrons

### v1

@@ Abstract
To illustrate the indirect detection constraints on the model parameters, we take the $e\mu$ flavor as an example. In Fig. 3, the solid curves represent the coupling strength required to achieve the correct relic density as a function of the DM mass, while the colored regions are excluded by indirect detection constraints. To obtain the relic density curves, we have used the velocity-averaged cross section formula developed in Gondolo:1990dk and matched it to the required value calculated in Steigman:2012nb . The dip corresponds to the point where $m_{\phi}\approx m_{\Phi}/2$, and is regulated by the estimated width $\Gamma_{\Phi}\sim 0.1\,m_{\Phi}/(4\pi)$. It can be seen that indirect detection constrains the DM mass to be above the GeV scale in this scenario. For the other two flavor scenarios ($e\tau$ and $\mu\tau$), the relic density curves remain applicable, but the lower bound of DM mass is approximately $m_{\tau}/2$ due to kinematic requirement. However, in these cases, the indirect constraints are weaker, indicating that the entire mass range is still viable. For the $e\mu$ case, our scenario can easily satisfy the constraints from the muonium oscillation, muon g2g-2, and forward-backward asymmetry of $e^{+}e^{-}\to\mu^{+}\mu^{-}$ derived in Endo:2020mev . Finally, the LFV processes $\mu\to e\gamma$ or $\mu\to 3e$ involve new LFC couplings, which can be adjusted to sufficiently small values to comply with experimental constraints.
@@ Acknowledgements
We thank Zihong Cheng for helpful discussions. This work was supported by Grants No. NSFC-12035008, No. NSFC-12305110, and No. NSFC-12347121.

### v2

@@ Abstract
To illustrate the indirect detection constraints on the model parameters, we take the $e\mu$ flavor as an example. In Fig. 3, the solid curves represent the coupling strength required to achieve the correct relic density as a function of the DM mass, while the colored regions are excluded by indirect detection constraints. To obtain the relic density curves, we have used the velocity-averaged cross section formula developed in Gondolo and Gelmini (1991) and matched it to the required value calculated in Steigman et al. (2012). The dip corresponds to the point where $m_{\phi}\approx m_{\Phi}/2$, and is regulated by the estimated width $\Gamma_{\Phi}\sim 0.1\,m_{\Phi}/(4\pi)$. It can be seen that indirect detection constrains the DM mass to be above the GeV scale in this scenario. For the other two flavor scenarios ($e\tau$ and $\mu\tau$), the relic density curves remain applicable, but the lower bound of DM mass is approximately $m_{\tau}/2$ due to kinematic requirement. However, in these cases, the indirect constraints are weaker, indicating that the entire mass range is still viable. For the $e\mu$ case, our scenario can easily satisfy the constraints from the muonium oscillation, muon g2g-2, and forward-backward asymmetry of $e^{+}e^{-}\to\mu^{+}\mu^{-}$ derived in Endo et al. (2020). Finally, the LFV processes $\mu\to e\gamma$ or $\mu\to 3e$ involve new LFC couplings, which can be adjusted to sufficiently small values to comply with experimental constraints.
@@ Acknowledgements
We thank the anonymous referee for several comments and suggestions that have helped us improve the presentation of our work. We thank Zihong Cheng for helpful discussions. This work was supported by Grants No. NSFC-12305110, No. NSFC-12035008, and No. NSFC-12347121.